OpenStudy (anonymous):

4(x-2) over 8=2(x+1) over 7??? help please

OpenStudy (anonymous):

help

OpenStudy (anonymous):

oh right this one multiply by the LCD 54 on both sides

OpenStudy (anonymous):

download Microsoft Mathematics if you can.

OpenStudy (anonymous):

56?

OpenStudy (anonymous):

I got 6... let me check

OpenStudy (anonymous):

\[\frac{4(x-2)}{8}=\frac{2(x+1)}{7}\] \[(7*8)\frac{4(x-2)}{8}=\frac{2(x+1)}{7}(7*8)\] 8 on the left cancels out, 7 on the right cancels out \[(7*4)(x-2)=(2*8)(x+1)\] can you solve it now?

OpenStudy (anonymous):

ok i got it thank yo i get confused math is not my favorite subject. i got three more if u got time

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

a plot of land is in h shape of a triangle. if one side is x meters, a second side is (5x-3) and a third side is (7x-5) meters express the perimeter of the lot as a simplified expresson in x

OpenStudy (anonymous):

the perimeter of the lot is__meters?

OpenStudy (anonymous):

perimeter is the sum of its sides

OpenStudy (anonymous):

x+(5x-3)+...

OpenStudy (anonymous):

then simplify like terms

OpenStudy (anonymous):

so would i distribute the x t each set of per?

OpenStudy (anonymous):

like x + 5x then x+3?

OpenStudy (anonymous):

13x-8 is what i got

OpenStudy (anonymous):

looks right

OpenStudy (anonymous):

the perimeter of a geometric figure is the sum of the lengths of its sides. the perimeter of the pentagon is 14 cm. a. write an equation for te perimeter b. solve the equation in part (a) c. find the length ofeach side

OpenStudy (anonymous):

i got part a

OpenStudy (anonymous):

how would i find the length of each side?

OpenStudy (anonymous):

no idea, i dont know the answer to a